OpenStudy (sidneyp):

(a + b)(a 2 - ab + b 2) I am trying to multiply polynomials by polynomials. Please help!!

jhonyy9 (jhonyy9):

so than i remember it right this wan being a formula of a^3 +b^3 (a+b)(a^2 -ab +b^2) = a^3 -a^2b +ab^2 +a^2b -ab^2 +b^3 = a^3 +b^3 hope helped

OpenStudy (triciaal):

you can shift the number 6 to get the symbol for exponents